in 1997, the two top selling cereals were general mills cheerios and kellogg's frosted flakes, with combined sales of 460,000,000. frosted flakes sales were 60,000,000 less then sales of cheerios, what were the 1997 sales far each brand
can anyone help

Answer by stanbon(75887)

in 1997, the two top selling cereals were general mills cheerios and kellogg's frosted flakes, with combined sales of 460,000,000. frosted flakes sales were 60,000,000 less then sales of cheerios, what were the 1997 sales far each brand
---------------
c + k = 460000000
k = c - 60000000
----------------------
Rewrite as follows:
c + k = 460000000
c - k = 60000000
-----------------
Add to solve for c:
2c = 520000000
c = 260000000 (cheerios)
--------
k = 460000000-260000000 = 200000000 (frosted flakes)
================
Cheers,
Stan H.